Hisako ran a cross-country race in which the course was divided into 3 equal legs. Hisako ran the first leg at 12 km/h and the last two legs at 16 km/h. The whole race took her 50 min. How long was each leg? How long was the entire course?

I know the solution is something like this:
12t + 16t + 16t = d
And I believe 50 minutes needs to be converted to hours, so 5/6 of an hour...but I'm not sure how to set up the equation so it's solvable. Please help.

Answer by checkley71(8403)

seeing as the distances are all the same you have three legs whose additive times equal 50 min.
d=rt or:
t=d/r
d/12+d/16+d/16=5/6 find a common denominator for the left hand side of this equation.
48 looks like a good choice.
(4d+3d+3d)/48=5/6
10d/48=5/6 cross multiply
60d=5*48
60d=240
d=240/60
d=4 k for the distance of each leg.
proof
4/12+4/16+4/16=5/6
(4*4+3*4+3*4)/48=5/6
(16+12+12)/48=5/6
40/48=5/6
5/6=5/6